Jarod is having a problem with rabbits getting into his vegetable garden, so he decides to fence it in. The length of the garden is 12 feet more than 3 times the width. He needs 80 feet of fencing to do this job. Find the length and width of the garden.
Found 2 solutions by killahzsmilez, richmusick:
Answer by killahzsmilez(5)

Let x equal the width.
Let 3x+12 equal the length.
x + 3x + 12 = 80

x + 3x + 12 = 80
4x + 12 = 80
- 12 -12
4x = 68
/4 /4
x = 17

Check
(17) + 3(17) + 12 = 80
(17) + (51) + 12 = 80
(17) + 63 = 80
80 = 80
The width is 17 feet and the length is 63 feet.

Answer by richmusick(3)

Let x equal the width. (w)
Let 3x+12 equal the length. (l)
P = 2w + 2l
2(x) + 2(3x + 12) = 80
2x + 6x + 24 = 80
8x + 24 = 80
- 24 -24
8x = 56
/8 /8
x = 7
3x + 12 = 21 + 12 = 33
Check
2(7) + 2(33) = 80
14 + 66 = 80
80 = 80
The width is 7 feet and the length is 33 feet.